M csu Blackboard L.. 0 Send Link D Save to MendeleyWatch Bulls vs Sp. Lngage385 Mail S Halliday, Fundamentals of Physics, 10e University Physics ractice Assignment Gradebook ORION Downloadable eTextbook gnment FULL SCREEN PRINTER VERSION Chapter 26, Problem 001 During the 4.3 min a 4.5 A current is set up in a wire, how many (a) coulombs and (b) electrons pass through any cross section of the wires width? (a) Number (b) Number Click if you would like to Show Work for this question: Units Units ( Open Show Work Question Attempts: Unlimited SAVE FOR LATERSUBMIT udy
Answer #1

Charge =current *time

Time should be in seconds and current in amperes

Charge=4 5*4.3*60=1161 C

(b) No of electrons =charge/charge of 1 electron

Charge of 1 e=1.6*10^(-19) C

So no of electron=7.26*10^(21)